Math 485: Graph Theory

Stephen G. Simpson 305 McAllister, 863-0775

                  simpson@math.psu.edu

August 24, 2009

This is an introductory graph theory course. Graph theory is a branch of mathematics which is valuable from the viewpoint of both theory and applications. It should be of interest to students of mathematics, computer science, engineering, and other fields.

The required textbook for this course is Introduction to Graph Theory, Second Edition, by Douglas B. West (Prentice Hall, 2001, ISBN 0-13-014400-2).

The web page for the course is

         http://www.math.psu.edu/simpson/courses/math485/.

We are meeting Mondays, Wednesdays, and Fridays, 11:15 AM to 12:05 PM, in 216 Boucke. I am usually in my office Monday through Friday, 9:00 AM to 5:00 PM. Office hours are by appointment, or you can drop in. Grades will be based on homework assignments, an in-class midterm exam, and a final exam.

Academic Integrity Statement

Academic dishonesty is not limited to simply cheating on an exam or assignment. The following is quoted directly from the PSU Faculty Senate Policies for Students regarding academic integrity and academic dishonesty: “Academic integrity is the pursuit of scholarly activity free from fraud and deception and is an educational objective of this institution. Academic dishonesty includes, but is not limited to, cheating, plagiarizing, fabricating of information or citations, facilitating acts of academic dishonesty by others, having unauthorized possession of examinations, submitting work of another person or work previously used without informing the instructor, or tampering with the academic work of other students.” All University and Eberly College of Science policies regarding academic integrity/academic dishonesty apply to this course and the students enrolled in this course. Refer to the following URL for further details on the academic integrity policies of the Eberly College of Science:

         http://www.science.psu.edu/academic/Integrity/index.html

Each student in this course is expected to work entirely on her/his own while taking any exam, to complete assignments on her/his own effort without the assistance of others unless directed otherwise by the instructor, and to abide by University and Eberly College of Science policies about academic integrity and academic dishonesty. Academic dishonesty can result in assignment of “F” by the course instructors or “XF” by Judicial Affairs as the final grade for the student.